Chastain Corporation is trying to determine the effect of its inventory turnover ratio and days sales outstanding (DSO) on its cash conversion cycle. Chastain's 2019 sales (all on credit) were $226,000, its cost of goods sold is 80% of sales, and it earned a net profit of 3%, or $6,780. It turned over its inventory 6 times during the year, and its DSO was 33 days. The firm had fixed assets totaling $30,000. Chastain's payables deferral period is 40 days. Assume 365 days in year for your calculations.

A. Calculate Chastain's cash conversion cycle. Do not round intermediate calculations. Round your answer to two decimal places.

days

B. Assuming Chastain holds negligible amounts of cash and marketable securities, calculate its total assets turnover and ROA. Do not round intermediate calculations. Round your answers to two decimal places.

Total assets turnover:

ROA: %

C. Suppose Chastain's managers believe that the inventory turnover can be raised to 8.6 times. What would Chastain's cash conversion cycle, total assets turnover, and ROA have been if the inventory turnover had been 8.6 for 2019? Do not round intermediate calculations. Round your answers to two decimal places.

Cash conversion cycle: days

Total assets turnover:

ROA: %
Answer #1

Answer a.

Days Inventory Outstanding = 365/ Inventory Turnover
Days Inventory Outstanding = 365 / 6
Days Inventory Outstanding = 60.83 days

Days Sales Outstanding = 33.00 days
Days Payable Outstanding = 40.00 days

Cash Conversion Cycle = Days Inventory Outstanding + Days Sales Outstanding – Days Payable Outstanding
Cash Conversion Cycle = 60.83 + 33.00 - 40.00
Cash Conversion Cycle = 53.83 days

Answer b.

Days Sales Outstanding = 365 * Accounts Receivable / Annual Sales
33.00 = 365 * Accounts Receivable / $226,000
Accounts Receivable = $20,432.88

Cost of Goods Sold = 80% * Sales
Cost of Goods Sold = 80% * $226,000
Cost of Goods Sold = $180,800

Inventory Turnover = Cost of Goods Sold / Inventories
6 = $180,800 / Inventories
Inventories = $30,133.33

Total Assets = Accounts Receivable + Inventories + Fixed Assets
Total Assets = $20,432.88 + $30,133.33 + $30,000.00
Total Assets = $80,566.21

Total Assets Turnover = Annual Sales / Total Assets
Total Assets Turnover = $226,000 / $80,566.21
Total Assets Turnover = 2.81

Return on Assets = Net Income / Total Assets
Return on Assets = $6,780.00 / $80,566.21
Return on Assets = 0.0842 or 8.42%

Answer c.

Days Inventory Outstanding = 365/ Inventory Turnover
Days Inventory Outstanding = 365 / 8.60
Days Inventory Outstanding = 42.44 days

Cash Conversion Cycle = Days Inventory Outstanding + Days Sales Outstanding – Days Payable Outstanding
Cash Conversion Cycle = 42.44 + 33.00 - 40.00
Cash Conversion Cycle = 35.44 days

Inventory Turnover = Cost of Goods Sold / Inventories
8.60 = $180,800 / Inventories
Inventories = $21,023.26

Total Assets = Accounts Receivable + Inventories + Fixed Assets
Total Assets = $20,432.88 + $21,023.26 + $30,000.00
Total Assets = $71,456.14

Total Assets Turnover = Annual Sales / Total Assets
Total Assets Turnover = $226,000 / $71,456.14
Total Assets Turnover = 3.16

Return on Assets = Net Income / Total Assets
Return on Assets = $6,780.00 / $71,456.14
Return on Assets = 0.0949 or 9.49%